HUBBLE SPACE TELESCOPE DAILY REPORT #1281 PERIOD COVERED: 0000Z Daily Status Report as of 342/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed WF/PC-2 5476 (Continuum Snapshots of 3CR Radio Galaxies) The WF/PC-2 was used to make snapshot images of radio galaxy 3C169.0 as part of a continuum survey and to study the galaxy's morphology. The observations were executed as scheduled, and no anomalies were noted. 1.2 Completed WF/PC-2 5764 (WF/PC-2 Cycle 4 Calibration: Internal Flats - Repeat) The WF/PC-2 was used to make internal flat calibration images using the VIS lamp and internal sources. The observations were executed as scheduled, and no anomalies were noted. 1.3 Completed FGS 5316 (Masses of Pre-Main Sequence Stars: Cycle 4 Medium) The Fine Guidance Sensors were used in the TRANS Mode to make astrometric observations of the pre-main sequence binary star ZZ-Tau in the Taurus star forming region. The FGS data will provide the apparent orbits. The observations were executed as scheduled, and no anomalies were noted. 1.4 Completed WF/PC-2 5370 (HST Medium-Deep Survey: Cycle 4 High, High Galactic Latitude) The WF/PC-2 was used in parallel with the prime science observations to make high galactic latitude observations as part of a medium-deep sky survey. The observations were completed as planned, and no problems were reported. 1.5 Completed GHRS/2 5769 (UV Spectroimaging of Saturn and Search for Exogenous Water in the Atmosphere) The High Resolution Spectrograph (Side Two) was used to make ultraviolet spectrographic observations of the North and South Hemispheres of the Planet Saturn to look for exogenous water. The target acquisition was performed on the moon Mimus, followed by offsets to several regions on Saturn. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Three Sets of FOC/96 5705 (A Color-Magnitude Array of NGC-7099: Cycle 4 Carry-over) The Faint Object Camera (f/96) was used to make a series of observations of the globular cluster NGC-7099 in two bands which will be used to build a color-magnitude array. Observations were made at the core of the cluster, and at offsets from the core. The observations were completed as planned, and no problems were reported 1.7 Completed Three Sets of WF/PC-2 5571 (WF/PC-2 Cycle 4 Calibration: Earthflats, Smaller Set) The WF/PC-2 was used to make several ultraviolet and narrow-band filter Earth flat calibration observations to correct for the telescope's illumination pattern. The observations were completed as planned, and no problems were reported. 1.8 Completed WF/PC-2 5479 (Subarcsecond Structure in Nearby AGNs) The WF/PC-2 was used to make snapshot observations of the nearby active galactic nuclei of NGC-6211 to study the interactions between the central engine, the stars, and the interstellar matter in the AGN. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed WF/PC-2 5397 (Determination of the Extragalactic Distance Scale: Cycle 4 High) The WF/PC-2 was used to make images of galaxy NGC-3351 to be used in measuring the periods of Cepheid variable stars within the galaxy. The observations were completed as planned, and no anomalies were noted. 1.10 Completed FOS/RD 5535 (Cycle 4 Internal Wavelength Calibration) The Faint Object Spectrograph was used to make internal calibration observations to monitor the stability of the wavelength scale for different gratings. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 7 Successful : 7 Scheduled Reacquisitions : 7 Successful : 7 2.2 FHST Updates: Scheduled : 17 Successful : 17 2.3 Significant HSTARs: None 2.4 Operations Notes: Initial STGT operations continued with the introduction of the East Satellite. There was only one TTR filed -- for a 2 minute, 55 second recoverable 32k data loss due to a problem with the integrated receiver. The Gyro #5 motor current transition from a nominal 98 mA to 156.8 mA had a signature very similar to the original Gyro #5 failure, with the exception that a current spike was not previously observed. Operational limits have been adjusted, and Gyro #5 data is being closely monitored. Operations were maintained, and there was no impact to science observations. 3.0 ANOMALY STATUS: 3.1 Significant New Anomalies: None 3.2 Significant Unresolved Anomalies: None 4.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH -END-